A hydrogen atom undergoes a transition form n=6 to n=2. Find the energy and frequency of the emitted photon
Answer #1

As energy of electron in hydrogen atom in n'th level is

13.6 En 2-

for n=2

  E_{n}=-\frac{13.6}{2^{2}}eV

for n=6

E_{n}=-\frac{13.6}{6^{2}}eV

energy of emitted photon is

  \Delta E=-\frac{13.6}{6^{2}}eV-(-\frac{13.6}{2^{2}}eV)=3.022eV

\Delta E=3.022eV

energy of emitted photon is 3.022eV.

frequency of emitted photon is

\nu =\frac{\Delta E}{h} where h is plank's constant =4.136 * 10-15eV-s

  \nu =\frac{3.022eV}{4.136\times 10^{-15}eV-s}=0.7306576\times 10^{15}

  \nu =0.7306576\times 10^{15}

frequency of photon emitted 0.7306576 * 1015 s-1.